OpenStudy (anonymous):

solve equations 36.) 1/3x+1/2=2 37.)7x=x-18 38.) 8r-36=-r 46.) 3(w-6)=-7(w+4) Does anyone know how to do equations w/fractions too????

thomaster (thomaster):

\(\large\dfrac{1}3x+\dfrac{1}2=2\) What's the first you wanna do?

OpenStudy (anonymous):

make the fraction bars whole numbers??

thomaster (thomaster):

Exactly, and how do we do that?

OpenStudy (anonymous):

idk??

thomaster (thomaster):

multiply the denominators of the 2 fractions. Then multiply everything in the equation with that.

OpenStudy (anonymous):

1/6x=2 ?????

thomaster (thomaster):

what are the denominators of the 2 fractions?

OpenStudy (anonymous):

3 and 2

thomaster (thomaster):

Yes, when you multiply those, you'll get 6. Now multiply the entire equation with 6

OpenStudy (anonymous):

1/6x=2 is what we have now right???

thomaster (thomaster):

No What is \(\large\dfrac{1}3x*6\ ?\)

OpenStudy (anonymous):

6/18?

thomaster (thomaster):

No... when you simplify 6/18, you'll get 1/3 When you multiply 1/3 with 3 what do you get?

OpenStudy (anonymous):

3 and 1/3

thomaster (thomaster):

No, we're multiplying here, not adding

thomaster (thomaster):

Imagine a cake, you cut it in 3 equal pieces. You take 1 piece, that's 1/3 piece. Multiply that with 3. You'll have 3/3 piece which is equal to 1 cake. So \(\large\dfrac{1}3x*3=1x\) but we're multiplying with 6, not 3. What will the answer be now?

OpenStudy (anonymous):

3/6

thomaster (thomaster):